# Stochastic Encodings for Active Feature Acquisition

### v1

@@ Appendix D Additional Experiment: Robustness to Noise
ACFlow 7.730±0.1397.730$\pm$ 0.1397.730 ± 0.139 7.754±0.1997.754$\pm$ 0.1997.754 ± 0.199 7.755±0.2757.755$\pm$ 0.2757.755 ± 0.275 8.255±0.1318.255$\pm$ 0.1318.255 ± 0.131
DIME 4.079±0.0644.079$\pm$ 0.0644.079 ± 0.064 4.688±0.2114.688$\pm$ 0.2114.688 ± 0.211 4.703±0.2764.703$\pm$ 0.2764.703 ± 0.276 5.368±0.2675.368$\pm$ 0.2675.368 ± 0.267
EDDI 9.197±0.2039.197$\pm$ 0.2039.197 ± 0.203 8.988±0.2858.988$\pm$ 0.2858.988 ± 0.285 9.216±0.1369.216$\pm$ 0.1369.216 ± 0.136 9.382±0.2459.382$\pm$ 0.2459.382 ± 0.245
Fixed MLP 6.009±0.0006.009$\pm$ 0.0006.009 ± 0.000 6.312±0.2026.312$\pm$ 0.2026.312 ± 0.202 7.321±0.3787.321$\pm$ 0.3787.321 ± 0.378 6.009±0.0006.009$\pm$ 0.0006.009 ± 0.000
GDFS 4.568±0.2194.568$\pm$ 0.2194.568 ± 0.219 4.566±0.1874.566$\pm$ 0.1874.566 ± 0.187 4.543±0.1644.543$\pm$ 0.1644.543 ± 0.164 5.583±0.2835.583$\pm$ 0.2835.583 ± 0.283
GSMRL 5.570±0.1275.570$\pm$ 0.1275.570 ± 0.127 5.495±0.1265.495$\pm$ 0.1265.495 ± 0.126 5.778±0.1205.778$\pm$ 0.1205.778 ± 0.120 6.858±0.2696.858$\pm$ 0.2696.858 ± 0.269
Opportunistic RL 4.201±0.0414.201$\pm$ 0.0414.201 ± 0.041 4.347±0.0914.347$\pm$ 0.0914.347 ± 0.091 4.720±0.1424.720$\pm$ 0.1424.720 ± 0.142 5.488±0.0845.488$\pm$ 0.0845.488 ± 0.084
Random 9.484±0.0069.484$\pm$ 0.0069.484 ± 0.006 9.495±0.0109.495$\pm$ 0.0109.495 ± 0.010 9.495±0.0109.495$\pm$ 0.0109.495 ± 0.010 9.495±0.0109.495$\pm$ 0.0109.495 ± 0.010
VAE 6.589±0.0886.589$\pm$ 0.0886.589 ± 0.088 6.866±0.0416.866$\pm$ 0.0416.866 ± 0.041 7.005±0.0867.005$\pm$ 0.0867.005 ± 0.086 7.095±0.0787.095$\pm$ 0.0787.095 ± 0.078
@@
ACFlow 7.527±0.2547.527$\pm$ 0.2547.527 ± 0.254 7.366±0.3257.366$\pm$ 0.3257.366 ± 0.325 7.927±0.3637.927$\pm$ 0.3637.927 ± 0.363 7.974±0.0727.974$\pm$ 0.0727.974 ± 0.072
DIME 4.581±0.2174.581$\pm$ 0.2174.581 ± 0.217 4.830±0.0574.830$\pm$ 0.0574.830 ± 0.057 5.093±0.2185.093$\pm$ 0.2185.093 ± 0.218 5.015±0.1245.015$\pm$ 0.1245.015 ± 0.124
EDDI 9.214±0.4159.214$\pm$ 0.4159.214 ± 0.415 9.550±0.3429.550$\pm$ 0.3429.550 ± 0.342 9.079±0.2449.079$\pm$ 0.2449.079 ± 0.244 9.571±0.2259.571$\pm$ 0.2259.571 ± 0.225
Fixed MLP 5.996±0.0005.996$\pm$ 0.0005.996 ± 0.000 6.693±0.3736.693$\pm$ 0.3736.693 ± 0.373 6.095±0.1006.095$\pm$ 0.1006.095 ± 0.100 6.494±0.3156.494$\pm$ 0.3156.494 ± 0.315
GDFS 4.484±0.1594.484$\pm$ 0.1594.484 ± 0.159 4.981±0.2114.981$\pm$ 0.2114.981 ± 0.211 5.655±0.3865.655$\pm$ 0.3865.655 ± 0.386 6.804±0.3846.804$\pm$ 0.3846.804 ± 0.384
GSMRL 6.227±0.1856.227$\pm$ 0.1856.227 ± 0.185 6.283±0.1466.283$\pm$ 0.1466.283 ± 0.146 6.525±0.1586.525$\pm$ 0.1586.525 ± 0.158 7.286±0.0887.286$\pm$ 0.0887.286 ± 0.088
Opportunistic RL 4.846±0.0214.846$\pm$ 0.0214.846 ± 0.021 5.042±0.0075.042$\pm$ 0.0075.042 ± 0.007 5.108±0.0645.108$\pm$ 0.0645.108 ± 0.064 5.246±0.0415.246$\pm$ 0.0415.246 ± 0.041
Random 9.499±0.0059.499$\pm$ 0.0059.499 ± 0.005 9.504±0.0079.504$\pm$ 0.0079.504 ± 0.007 9.504±0.0079.504$\pm$ 0.0079.504 ± 0.007 9.504±0.0079.504$\pm$ 0.0079.504 ± 0.007
VAE 6.667±0.1406.667$\pm$ 0.1406.667 ± 0.140 6.874±0.1616.874$\pm$ 0.1616.874 ± 0.161 7.057±0.1347.057$\pm$ 0.1347.057 ± 0.134 7.204±0.0947.204$\pm$ 0.0947.204 ± 0.094
@@
ACFlow 9.194±0.2789.194$\pm$ 0.2789.194 ± 0.278 9.105±0.2189.105$\pm$ 0.2189.105 ± 0.218 9.184±0.2489.184$\pm$ 0.2489.184 ± 0.248 9.284±0.2319.284$\pm$ 0.2319.284 ± 0.231
DIME 5.667±0.0385.667$\pm$ 0.0385.667 ± 0.038 5.771±0.2225.771$\pm$ 0.2225.771 ± 0.222 6.180±0.1616.180$\pm$ 0.1616.180 ± 0.161 6.213±0.1806.213$\pm$ 0.1806.213 ± 0.180
EDDI 9.794±0.1869.794$\pm$ 0.1869.794 ± 0.186 9.969±0.1569.969$\pm$ 0.1569.969 ± 0.156 9.906±0.2129.906$\pm$ 0.2129.906 ± 0.212 9.710±0.1369.710$\pm$ 0.1369.710 ± 0.136
Fixed MLP 7.999±0.0007.999$\pm$ 0.0007.999 ± 0.000 7.999±0.0007.999$\pm$ 0.0007.999 ± 0.000 7.999±0.0007.999$\pm$ 0.0007.999 ± 0.000 7.999±0.0007.999$\pm$ 0.0007.999 ± 0.000
GDFS 5.587±0.2015.587$\pm$ 0.2015.587 ± 0.201 6.839±0.3386.839$\pm$ 0.3386.839 ± 0.338 7.179±0.1347.179$\pm$ 0.1347.179 ± 0.134 7.814±0.3497.814$\pm$ 0.3497.814 ± 0.349
GSMRL 8.199±0.0678.199$\pm$ 0.0678.199 ± 0.067 8.326±0.2338.326$\pm$ 0.2338.326 ± 0.233 8.347±0.1228.347$\pm$ 0.1228.347 ± 0.122 8.627±0.1868.627$\pm$ 0.1868.627 ± 0.186
Opportunistic RL 5.850±0.0725.850$\pm$ 0.0725.850 ± 0.072 5.859±0.0635.859$\pm$ 0.0635.859 ± 0.063 6.135±0.1146.135$\pm$ 0.1146.135 ± 0.114 6.672±0.1056.672$\pm$ 0.1056.672 ± 0.105
Random 9.987±0.0089.987$\pm$ 0.0089.987 ± 0.008 9.991±0.0069.991$\pm$ 0.0069.991 ± 0.006 9.991±0.0069.991$\pm$ 0.0069.991 ± 0.006 9.991±0.0069.991$\pm$ 0.0069.991 ± 0.006
VAE 7.888±0.0657.888$\pm$ 0.0657.888 ± 0.065 7.968±0.0687.968$\pm$ 0.0687.968 ± 0.068 8.042±0.1028.042$\pm$ 0.1028.042 ± 0.102 8.469±0.0648.469$\pm$ 0.0648.469 ± 0.064
@@ Appendix F Indicator Example
[…]
@@
[…]
@@
[…]
If we consider the first $d$ features, we can again apply symmetry to calculate this quantity for one feature - feature 111 - and apply it to all of them. In $\frac{d-1}{d}$ cases the entropy is zero, since we will have all of the information required. However if $x_{d+1}=1$, then $H(Y|\mathbf{x}_{U})=\log 2$, since we don’t know the value of feature 111 and therefore $Y$ has equal likelihood of being 0 or 1, this happens in $\frac{1}{d}$ cases so this quantity is $\frac{\log 2}{d}$ for the first $d$ features.
For the indicator, $p(Y=0|\mathbf{x}_{U})$ is the proportion of the first $d$ features that are 0 for a given sample $\mathbf{x}_{U}$. All features are independent with probability 0.50.50.5 of being 0, so the expression is given by a binomial distribution with $d$ trials
@@
It is not immediately clear that this is larger than the quantity $\frac{\log 2}{d}$ for the other features. The first thing we can do is calculate this quantity when $d=3$, which gives 0.4770.4770.477, and this is larger than $\frac{\log 2}{3}=0.231$. And the next thing is to notice that this quantity is increasing with $d$, since as $d$ gets larger there will be more probability mass at $k=\frac{d}{2}$. As $d\xrightarrow{}\infty$ the binomial distribution becomes Gaussian with mean $\frac{d}{2}$ and variance $\frac{d}{4}$, so $\frac{k}{d}$ will approximately be distributed normally with mean $\frac{1}{2}$ and standard deviation $\frac{1}{2\sqrt{d}}$. Therefore this quantity asymptotes towards $\log 2$.
Therefore for $d\geq 3$, this objective will choose the indicator first, and not the other features (for $d=2$ all features are scored the same, and for $d=1$ the indicator is not the optimal choice). After choosing the indicator, the second selection is trivial. The relevant feature has non-zero CMI, all other features are independent of the label conditioned on the indicator so they have zero CMI. Therefore the correct feature is chosen. This strategy’s expected number of acquisitions is 222, which is less than $3-\frac{1}{d}$.
@@ Appendix G Entropy Example
Consider a feature vector with two features, $X\in\{1,2,3\}^{2}$, and a label with three possible classes, $Y\in\{1,2,3\}$. To generate the label from the features, we start with a vector with three zeros $\mathbf{v}=[0.0,0.0,0.0]$. The value of $x_{1}$ tells us which value in $\mathbf{v}$ to reduce by 888. And the value of $x_{2}$ tells us which value in $\mathbf{v}$ to increase by 6.56.56.5. For example if $\mathbf{x}=[1,2]$, then $\mathbf{v}=[-8.0,6.5,0.0]$. $p(y|\mathbf{x})$ is then given by the softmax of $\mathbf{v}$. With these rules, we can calculate all possible feature values and probabilities in Table 5.
@@
Missing 111 0.699120.699120.69912 0.150440.150440.15044 0.150440.150440.15044 0.82016 0.27845
Missing 222 0.150440.150440.15044 0.699120.699120.69912 0.150440.150440.15044 0.82016 0.27845
Missing 333 0.150440.150440.15044 0.150440.150440.15044 0.699120.699120.69912 0.82016 0.27845
Feature 2 is consistently more useful for identifying the most likely class, since it can always identify a class with likelihood 0.6990.6990.699, and feature 1 is better at identifying which class is least likely. However, feature 1 consistently has a lower entropy, so it would be selected before feature 2 by CMI.
@@ Appendix H Dataset Details
  • Class 1: Features 1, 2, 3 have mean [0,0,0][0,0,0][ 0 , 0 , 0 ] and diagonal standard deviation [0.1,0.1,0.1][0.1,0.1,0.1][ 0.1 , 0.1 , 0.1 ].
  • Class 2: Features 2, 3, 4 have mean [1,0,0][1,0,0][ 1 , 0 , 0 ]and diagonal standard deviation [0.1,0.1,0.1][0.1,0.1,0.1][ 0.1 , 0.1 , 0.1 ].
  • Class 3: Features 3, 4, 5 have mean [0,1,0][0,1,0][ 0 , 1 , 0 ]and diagonal standard deviation [0.1,0.1,0.1][0.1,0.1,0.1][ 0.1 , 0.1 , 0.1 ].
  • Class 4: Features 4, 5, 6 have mean [1,1,0][1,1,0][ 1 , 1 , 0 ]and diagonal standard deviation [0.1,0.1,0.1][0.1,0.1,0.1][ 0.1 , 0.1 , 0.1 ].
  • Class 5: Features 5, 6, 7 have mean [0,0,1][0,0,1][ 0 , 0 , 1 ]and diagonal standard deviation [0.1,0.1,0.1][0.1,0.1,0.1][ 0.1 , 0.1 , 0.1 ].
  • Class 6: Features 6, 7, 8 have mean [1,0,1][1,0,1][ 1 , 0 , 1 ] and diagonal standard deviation [0.1,0.1,0.1][0.1,0.1,0.1][ 0.1 , 0.1 , 0.1 ].
  • Class 7: Features 7, 8, 9 have mean [0,1,1][0,1,1][ 0 , 1 , 1 ] and diagonal standard deviation [0.1,0.1,0.1][0.1,0.1,0.1][ 0.1 , 0.1 , 0.1 ].
  • Class 8: Features 8, 9, 10 have mean [1,1,1][1,1,1][ 1 , 1 , 1 ] and diagonal standard deviation [0.1,0.1,0.1][0.1,0.1,0.1][ 0.1 , 0.1 , 0.1 ].
@@ I.1 General Model Details
Input Representation.   Continuous features are represented as $[\mathbf{x}\odot\mathbf{m},\mathbf{m}]$. Categorical features use a one-hot encoding, where we include an additional class to indicate a missing feature. Continuous and categorical representations are then concatenated as input to the main model.
Deep Networks.   All deep networks are given by MLPs with ReLU activation followed by Batch Normalization (Ioffe & Szegedy, 2015). All hidden layers in a given network are the same width, which is a hyperparameter that can be tuned as well as the number of hidden layers. The exception to this is the Opportunistic RL model, where we replace Batch Normalization with dropout with 0.50.50.5 probability in accordance with the method’s implementation (Kachuee et al., 2019a).

### v3

@@ Appendix D Additional Experiment: Robustness to Noise
[…]
DIME $4.079\pm 0.064$ $4.688\pm 0.211$ $4.703\pm 0.276$ $5.368\pm 0.267$
EDDI $9.197\pm 0.203$ $8.988\pm 0.285$ $9.216\pm 0.136$ $9.382\pm 0.245$
Fixed MLP $6.009\pm 0.000$ $6.312\pm 0.202$ $7.321\pm 0.378$ $6.009\pm 0.000$
GDFS $4.568\pm 0.219$ $4.566\pm 0.187$ $4.543\pm 0.164$ $5.583\pm 0.283$
GSMRL $5.570\pm 0.127$ $5.495\pm 0.126$ $5.778\pm 0.120$ $6.858\pm 0.269$
Opportunistic RL $4.201\pm 0.041$ $4.347\pm 0.091$ $4.720\pm 0.142$ $5.488\pm 0.084$
Random $9.484\pm 0.006$ $9.495\pm 0.010$ $9.495\pm 0.010$ $9.495\pm 0.010$
VAE $6.589\pm 0.088$ $6.866\pm 0.041$ $7.005\pm 0.086$ $7.095\pm 0.078$
@@
ACFlow $7.527\pm 0.254$ $7.366\pm 0.325$ $7.927\pm 0.363$ $7.974\pm 0.072$
DIME $4.581\pm 0.217$ $4.830\pm 0.057$ $5.093\pm 0.218$ $5.015\pm 0.124$
EDDI $9.214\pm 0.415$ $9.550\pm 0.342$ $9.079\pm 0.244$ $9.571\pm 0.225$
Fixed MLP $5.996\pm 0.000$ $6.693\pm 0.373$ $6.095\pm 0.100$ $6.494\pm 0.315$
GDFS $4.484\pm 0.159$ $4.981\pm 0.211$ $5.655\pm 0.386$ $6.804\pm 0.384$
GSMRL $6.227\pm 0.185$ $6.283\pm 0.146$ $6.525\pm 0.158$ $7.286\pm 0.088$
Opportunistic RL $4.846\pm 0.021$ $5.042\pm 0.007$ $5.108\pm 0.064$ $5.246\pm 0.041$
Random $9.499\pm 0.005$ $9.504\pm 0.007$ $9.504\pm 0.007$ $9.504\pm 0.007$
VAE $6.667\pm 0.140$ $6.874\pm 0.161$ $7.057\pm 0.134$ $7.204\pm 0.094$
@@
ACFlow $9.194\pm 0.278$ $9.105\pm 0.218$ $9.184\pm 0.248$ $9.284\pm 0.231$
DIME $5.667\pm 0.038$ $5.771\pm 0.222$ $6.180\pm 0.161$ $6.213\pm 0.180$
EDDI $794\pm 0.186 9.9$$69\pm 0.156 9.9$$906\pm 0.212$ $9.710\pm 0.136$
Fixed MLP $7.999\pm 0.000$ $7.999\pm 0.000$ $7.999\pm 0.000$ $7.999\pm 0.000$
GDFS $5.587\pm 0.201$ $6.839\pm 0.338$ $7.179\pm 0.134$ $7.814\pm 0.349$
GSMRL $8.199\pm 0.067$ $8.326\pm 0.233$ $8.347\pm 0.122$ $8.627\pm 0.186$
Opportunistic RL $5.850\pm 0.072$ $5.859\pm 0.063$ $135\pm 0.114 6.6$$672\pm 0.105$
Random $987\pm 0.008 9.9$$91\pm 0.006 9.9$$91\pm 0.006 9.9$$991\pm 0.006$
VAE $7.888\pm 0.065$ $7.968\pm 0.068$ $8.042\pm 0.102$ $8.469\pm 0.064$
@@ Appendix F Indicator Example
$I(Y;X_{i})=H(Y)-\int H(Y|x_{i})p(x_{i})dx_{i}.$
The entropy of the label is $\log 2$ since there is equal chance of being 0 or 1. Again, using the symmetry of the system, the entropy of $Y$ if $X_{i}=0$ is the same as if $X_{i}=1$, so we only calculate for one case. When $X_{i}=0$, the probability of $Y=0$ is $\frac{1}{d}\times 1+\frac{d-1}{d}\times\frac{1}{2}$, since in $\frac{1}{d}$ cases it takes the exact value of $X_{i}$ based on the value of the indicator, and in $\frac{d-1}{d}$ cases $Y$ is given by a different unknown feature value. This gives $p(Y=0|X_{i}=0)=\frac{d+1}{2d}$. The expression for binary entropy, $p\log(p)-(1-p)\log(1-p)$ is maximized by $p=0.5$, giving $\log 2$. Since $p(Y=0|X_{i}=0)>0.5$, the entropy is lower than $\log 2$ in this case. Exploiting the symmetry of the system we conclude that $\int H(Y|x_{i})p(x_{i})dx_{i}<\log 2$, and therefore $I(Y;X_{i})>0$.
@@
Initially there are no features, so the acquisition objective is $\int I(X_{i};Y|\mathbf{x}_{U})p(\mathbf{x}_{U})d\mathbf{x}_{U}$. Writing this in terms of entropies gives
$\int I(X_{i};Y|\mathbf{x}_{U})p(\mathbf{x}_{U})d\mathbf{x}_{U}=\int\Bigg{(}H(Y|\mathbf{x}_{U})-\int H(Y|x_{i},\mathbf{x}_{U})p(x_{i}|\mathbf{x}_{U})dx_{i}\Bigg{)}p(\mathbf{x}_{U})d\mathbf{x}_{U}.$
@@
$\int H(Y|\mathbf{x}_{U})p(\mathbf{x}_{U})d\mathbf{x}_{U}.$
If we consider the first $d$ features, we can again apply symmetry to calculate this quantity for one feature - feature $1$ - and apply it to all of them. In $\frac{d-1}{d}$ cases the entropy is zero, since we will have all of the information required. However if $x_{d+1}=1$, then $H(Y|\mathbf{x}_{U})=\log 2$, since we don’t know the value of feature $1$ and therefore $Y$ has equal likelihood of being 0 or 1, this happens in $\frac{1}{d}$ cases so this quantity is $\frac{\log 2}{d}$ for the first $d$ features.
For the indicator, $p(Y=0|\mathbf{x}_{U})$ is the proportion of the first $d$ features that are 0 for a given sample $\mathbf{x}_{U}$. All features are independent with probability $0.5$ of being 0, so the expression is given by a binomial distribution with $d$ trials
@@
It is not immediately clear that this is larger than the quantity $\frac{\log 2}{d}$ for the other features. The first thing we can do is calculate this quantity when $d=3$, which gives $0.477$, and this is larger than $\frac{\log 2}{3}=0.231$. And the next thing is to notice that this quantity is increasing with $d$, since as $d$ gets larger there will be more probability mass at $k=\frac{d}{2}$. As $d\xrightarrow{}\infty$ the binomial distribution becomes Gaussian with mean $\frac{d}{2}$ and variance $\frac{d}{4}$, so $\frac{k}{d}$ will approximately be distributed normally with mean $\frac{1}{2}$ and standard deviation $\frac{1}{2\sqrt{d}}$. Therefore this quantity asymptotes towards $\log 2$.
Therefore for $d\geq 3$, this objective will choose the indicator first, and not the other features (for $d=2$ all features are scored the same, and for $d=1$ the indicator is not the optimal choice). After choosing the indicator, the second selection is trivial. The relevant feature has non-zero CMI, all other features are independent of the label conditioned on the indicator so they have zero CMI. Therefore the correct feature is chosen. This strategy’s expected number of acquisitions is $2$, which is less than $3-\frac{1}{d}$.
@@ Appendix G Entropy Example
Consider a feature vector with two features, $X\in\{1,2,3\}^{2}$, and a label with three possible classes, $Y\in\{1,2,3\}$. To generate the label from the features, we start with a vector with three zeros $\mathbf{v}=[0.0,0.0,0.0]$. The value of $x_{1}$ tells us which value in $\mathbf{v}$ to reduce by $8$. And the value of $x_{2}$ tells us which value in $\mathbf{v}$ to increase by $6.5$. For example if $\mathbf{x}=[1,2]$, then $\mathbf{v}=[-8.0,6.5,0.0]$. $p(y|\mathbf{x})$ is then given by the softmax of $\mathbf{v}$. With these rules, we can calculate all possible feature values and probabilities in Table 5.
@@
Missing 11 $0.69912$ $0.15044$ $0.15044$ 0.82016 0.27845
Missing 22 $0.15044$ $0.69912$ $0.15044$ 0.82016 0.27845
Missing 33 $0.15044$ $0.15044$ $0.69912$ 0.82016 0.27845
Feature 2 is consistently more useful for identifying the most likely class, since it can always identify a class with likelihood $0.699$, and feature 1 is better at identifying which class is least likely. However, feature 1 consistently has a lower entropy, so it would be selected before feature 2 by CMI.
@@ Appendix H Dataset Details
  • Class 1: Features 1, 2, 3 have mean $[0,0,0]$ and diagonal standard deviation $[0.1,0.1,0.1]$.
  • Class 2: Features 2, 3, 4 have mean [1,0,0][1,0,0]and diagonal standard deviation $[0.1,0.1,0.1]$.
  • Class 3: Features 3, 4, 5 have mean [0,1,0][0,1,0]and diagonal standard deviation $[0.1,0.1,0.1]$.
  • Class 4: Features 4, 5, 6 have mean [1,1,0][1,1,0]and diagonal standard deviation $[0.1,0.1,0.1]$.
  • Class 5: Features 5, 6, 7 have mean [0,0,1][0,0,1]and diagonal standard deviation $[0.1,0.1,0.1]$.
  • Class 6: Features 6, 7, 8 have mean $[1,0,1]$ and diagonal standard deviation $[0.1,0.1,0.1]$.
  • Class 7: Features 7, 8, 9 have mean $[0,1,1]$ and diagonal standard deviation $[0.1,0.1,0.1]$.
  • Class 8: Features 8, 9, 10 have mean $[1,1,1]$ and diagonal standard deviation $[0.1,0.1,0.1]$.
@@ I.1 General Model Details
Input Representation.   Continuous features are represented as $\mathbf{x}\odot\mathbf{m},\mathbf{m}]$. Categorical features use a one-hot encoding, where we include an additional class to indicate a missing feature. Continuous and categorical representations are then concatenated as input to the main model.
Deep Networks.   All deep networks are given by MLPs with ReLU activation followed by Batch Normalization (Ioffe & Szegedy, 2015). All hidden layers in a given network are the same width, which is a hyperparameter that can be tuned as well as the number of hidden layers. The exception to this is the Opportunistic RL model, where we replace Batch Normalization with dropout with $0.5$ probability in accordance with the method’s implementation (Kachuee et al., 2019a).